The local movie theater sold 100 tickets for $780. If children's tickets cost $6 and adult tickets $12, how many of each were sold?
Answer by Maths68(1474)

Let
Children's ticket = x
Adult's ticket = y
Cost of Children's tickets = 6x
Cost of Adult's tickets = 12y
Children's ticket + Adult's ticket = Total Tickets sold
x+y=100
x=100-y ..............(1)
cost of children tickets + cost of Adult's ticket = cost of total ticket sold
6x+12y=780...........(2)
Put the value of x from (1) to (2)
6x+12y=780
6(100-y)+12y=780
600-6y+12y=780
600+6y=780
6y=780-600
6y=180
y=180/6
y=30
put the value of y in (1)
x=100-y
x=100-30
x=70
Children's tickets sold = 70
Adult's tickets sold = 30